OpenStudy (anonymous):

Given a = -3, b = 4 and c = -5, evaluate -|a - b| + c.

OpenStudy (anonymous):

-|-3 - 4| + (-5) Scratch the ||'s and get -7+(-5) which equals what?

OpenStudy (imstuck):

Filling in your values you get-3 - 4 = -7 but the absolute value of -7 is 7. NOW take the negative of it because of the negative outside the a.v. symbols. -7 - 5 = -12.

OpenStudy (anonymous):

I'm math dumb, lol. Thank you!

OpenStudy (anonymous):

Absolute Values are super easy, you'll see once you get into higher math levels.

OpenStudy (anonymous):

Can you help me with this one also. A=H/2(a+b); for H

OpenStudy (anonymous):

Use Mathway.com it literally helps me on most of my math.

OpenStudy (imstuck):

In this case, you would multiply both sides by 2(a + b) to get rid of it in the denominator. You would have 2A(a + b) = H